OpenStudy (anonymous):

The least common multiple of 20, 24, and 45 is _____. a) 30 b) 180 c) 360 d) 21,600

OpenStudy (compassionate):

Neither.

Parth (parthkohli):

\[20 = 2 \cdot 2 \cdot 5\]\[24 = 2 \cdot 2\cdot 2 \cdot 3\]\[45 = 3\cdot 3\cdot5\]So, \(\rm lcm(20,24,45) = 2\cdot2\cdot 3 \cdot 5 \cdot 2 \cdot 3\)

Parth (parthkohli):

@OneKoi: Hello?

OpenStudy (anonymous):

360

Parth (parthkohli):

If a simpler way is needed,\[\begin{array}{l|r} 5 & {20,24,45} \\ \hline 4 & 4,24,9 \\ \hline 3& 1,6,9 \\ \hline2 & 1,2,3 \\ \hline 3 & 1,1,3 \\\hline & 1,1,1 \end{array}\]

Parth (parthkohli):

So, \(5 \cdot 4 \cdot 3 \cdot 2 \cdot 3 = \cdots \)

hartnn (hartnn):

just an alternative method... list out multiples for all of them.. like 20---->20,40,60,80...... 24---->24,48,.... 45---->.... and see which multiple comes out to be common first. but its the crude(and easy) way to find, but long.

Parth (parthkohli):

@OneKoi: Do you know any of the three methods listed?

OpenStudy (anonymous):

@OneKoi is not in the room or in front of Computer I think...

Parth (parthkohli):

Seems so...